OpenStudy (anonymous):

A $145 scooter is on sale at 20% off. Write a proportion you could use to find the amount discounted (how much money you saved). How much will the scooter cost after the 20% is Help please taken off of the original price. Show your work for credit.

OpenStudy (anonymous):

145 * 0.2= 29 145-29=116 145-116= you save 29 dollars